OpenStudy (bacon):

5x+6=2x=15 How do i solve it?????!!!!

OpenStudy (anonymous):

I'm confused...is there really two equal signs?

OpenStudy (bacon):

oh oops sorry let me rewrite that

OpenStudy (bacon):

5x+6=2x+15

OpenStudy (bacon):

there

OpenStudy (anonymous):

Ok so what you want to do is move all the same variables on one side: 5x+6=2x+15 Subtract 2x and subtract 6 from both sides so both sides have the same variable: 3x=9 Now divide 3: x=3

OpenStudy (anonymous):

Got it?

OpenStudy (bacon):

yes

OpenStudy (bacon):

thank you so much!!!!

OpenStudy (bacon):

so would 7x-4=20+15 with the x=6?

OpenStudy (anonymous):

I think you're missing another x?

OpenStudy (bacon):

oh yup

OpenStudy (bacon):

sorry i keep on messing up...-_-

OpenStudy (bacon):

7x-4=20+3x

OpenStudy (bacon):

x=6

OpenStudy (anonymous):

Perfect! Great job! You can check if your answer is correct or not by plugging in the answer to x :)

OpenStudy (bacon):

thanks

OpenStudy (bacon):

this is kind of off topic but can you help me with this.... I need to find another example of what a dream deferred is like for this poem...any ideas? What happens to a dream deferred? Does it dry up like a raisin in the sun? Or fester like a sore-- And then run? Does it stink like rotten meat? Or crust and sugar over-- like a syrupy sweet? Maybe it just sags like a heavy load. Or does it explode? by Langston Hughes

OpenStudy (anonymous):

A dream differed is like fine port wine aging in a cask. You always want to break it out an enjoy it with your friends. If you leave it sit in the cask too long it tuns to vinegar and is spoiled for all.